7(5)n-1 Find the exact sum (no decimal approximations) of the series 2 - n=1 e series -. (no explanation needed) . 8
Answer #1

Solution The Series is th e ones n= 1 1-8 It is a infinite Gip with a = 5 = 1 / 2 and n = 5 The formula of sum is a so, s.. .